# Inferring Change Points in High-Dimensional Regression via Approximate Message Passing

### v2

@@ 4.2 Rectified Linear Regression with Change Points
For the rectified linear model in (4) with a single change point at $(0.6n)$, Figure 5 (right) plots the Hausdorff distance divided by $n$ as a function of $\delta$. We let $p=600$, $L=3$, ${\mathbb{P}}_{\bar{\bm{B}}}={\sf N}({\bm{0}},\delta\bm{I})$, $\sigma=0.3$, and average over ten trials. Similarly, Table 1 displays the performance of AMP in estimating the number of change points in the same setup, where we vary $\sigma$ from $0.5$ to $0.3$. We observe a close match between the AMP performance and that dictated by theory in Proposition 3. Figure 2 (right) plots the estimated posterior $p_{\bar{{\bm{\Psi}}}|\bm{V}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}(\cdot|\bm{\Theta}^{t},\bm{y})$ against its the limiting posterior $p_{\bar{{\bm{\Psi}}}|\bm{V}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}(\cdot|{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},{\bm{\Psi}},\bm{\varepsilon}))$, for the rectified linear model with $p=600,\sigma=0.5$, true change point locations at $n/3$ and $158n/$ 15, and $L=3$ with a uniform prior over all configurations with two change points. We empirically observe a match between the estimated and limiting posteriors as per Proposition 4.
@@ 4.3 Logistic Regression with Change Points
For the logistic model in (3) with change points at $0.35n$ and $0.7n$, Figure 5 (left) plots the Hausdorff distance normalized by $n$ for varying $\delta$. We choose $p=600$, $\Delta=n/5$, and $L=3$ with a uniform prior over all valid configurations. We let ${\mathbb{P}}_{\bar{\bm{B}}}={\sf N}\left({\bm{0}},\begin{bmatrix}15&0.2&0\\ 0.2&15&0\\ 0&0&15\end{bmatrix}\right)$. We observe a close match between the AMP performance and that dictated by the state evolution theory in (31).
@@
Figure 2 (middle) shows the estimated posterior $p(\cdot|\bm{\Theta}^{t},\bm{y})$ from (34) and the limiting (exact) posterior $p(\cdot|{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},{\bm{\Psi}},\bm{\varepsilon}))$ in the logistic model with two true change points at $n/3$ and $158n/$ 15. We let $p=600$, ${\mathbb{P}}_{\bar{\bm{B}}}={\sf N}({\bm{0}},20\bm{I})$, and subsample the domain $[0,1]$ by a factor of eight. We let $L=3$ and place uniform prior over all possible two-changepoint configurations. We observe a close match with the theory in Proposition 4, and concentration around the true change point locations as $\delta$ increases.
Figure 6 compares AMP with the binary segmentation algorithm proposed in Wang et al. (2023) for sparse logistic regression. We use $p=300,\Delta=0.15n$, and a sparse Bernoulli-Gaussian signal prior ${\mathbb{P}}_{\bar{\bm{B}}}=0.5{\sf N}({\bm{0}},15\delta\bm{I})+0.5\delta_{\bm{0}}$. The true change points are at $n/3$ and $158n/$ 15 with $L^{*}=3$. The hyperparameters for binary segmentation are chosen as suggested in the source code of Wang et al. (2023). Both algorithms can take advantage of prior information on the number of change points. In Fig. 6(a), both algorithms assume a maximum of $L=3$ or 4 signals while $L^{*}=3$. The algorithms are observed to achieve similar Hausdorff distance, with AMP enjoying a much lower runtime when assuming $L=3$. Fig. 6(b) compares the two algorithms by assuming there are exactly 3 signals. The performance of AMP improves significantly with larger $\delta$, both in terms of Hausdorff distance and runtime. We recall that the algorithm in Wang et al. (2023) is tailored to sparse logistic regression, whereas AMP can account for a broader range of signal priors.
@@
(a) Algorithms assume a maximum of $3$ or $4$ signals (i.e., $L=3$ or 4).
@@
We consider the Myocardial Infarction (MI) complications dataset from Golovenkin et al. (2020), which contains the medical information of $n=1700$ patients (samples) with MI complications. Each sample has $p=111$ medical features such as age, sex, heredity, and the presence of diabetes. The dataset also contains $12$ binary response variables for each patient relating to the state of the patient’s overall heart health, indicating the presence of complications such as ‘Atrial Fibration’ and ‘Chronic Heart Failure’ (CHF). We investigate the relation between the binary CHF response variable and the features of each patient, using the logistic model. We sort the data by age in the range $26-92$ and exclude age as a feature, with the aim of determining whether the relationship between the response and the features changes markedly at a certain age. We whiten the age-ordered $n\times(p-1)$ feature matrix, replacing missing values with interpolated estimates. We set $L=2,\Delta=n/500$, and select the signal prior ${\mathbb{P}}_{\bar{\bm{B}}}={\sf N}\left({\bm{0}},\begin{bmatrix}1&3/4\\ 3/4&1\end{bmatrix}\right)$ independently of the data. For $\pi_{\bar{{\bm{\Psi}}}}$ we first specify a uniform prior over the number of change points (0 or 1 change points with probability $\frac{1}{2}$ each); we then define a uniform prior on the change point location for the case of 1 change point. We compute the estimated posterior $p({\bm{\psi}}|\bm{\Theta}^{t},\bm{y})$ from (34), reporting the existence of a change point with posterior probability close to 1. Assuming one change point, we plot the estimated posterior over all possible single change point locations in Figure 7 (left). The plot displays strong concentration of the posterior within the age range $55-75$, attaining its maximum value at age $66$. This indicates the existence of a change point in the age range $55-75$ in how the presence of CHF relates to various medical features such as heredity, sex, and diabetes. This is consistent with findings in the medical literature, where older patients with the presence of CHF (aged $>75$ years) were found to be more often female, have fewer cardiovascular diseases, and have fewer associated risk factors than patients with CHF aged $55$ years or less (Stein et al., 2012; Azad and Lemay, 2014; Tromp et al., 2021).
@@ Appendix A Proof of State Evolution
We extend this lemma by considering an independent random matrix ${\bm{\Xi}}\in{\mathbb{R}}^{N\times L_{{\bm{\Xi}}}}$, serving as input to a set of pseudo-Lipschitz functions $\varphi:\mathbb{R}^{N\times(Q+L_{{\bm{\Xi}}})}\to\mathbb{R}^{N}$. We then analyze a similar AMP iteration whose denoising function $\tilde{f}^{t}$ takes $\varphi(\bm{A}\bm{W}_{0},{\bm{\Xi}})$ as input instead of $\varphi(\bm{A}\bm{W}_{0})$, initialized with an independent initializer $\mathbf{X}^{0}\in\mathbb{R}^{N\times Q}$:
@@
    $\bm{A}\in\mathbb{R}^{N\times N}$ is a GOE($N$) matrix, i.e., $\bm{A}=\mathbf{G}+\mathbf{G}^{\top}$ for $\mathbf{G}\in\mathbb{R}^{N\times N}$ with i.i.d. entries $G_{ij}\sim{\sf N}(0,1/(2N))$.
@@
    where $(\bm{Z}^{s},\bm{Z}^{t})\in(\mathbb{R}^{N\times Q})^{2}$,$(\bm{Z}^{s},\bm{Z}^{t})\sim{\sf N}(0,\bm{\kappa}\otimes\mathbf{I}_{N})$.
@@
$\Phi_{N}\left(\mathbf{X}^{0},\mathbf{X}^{1},...,\mathbf{X}^{t},\bm{A}\bm{W}_{0};\bm{W}_{0},{\bm{\Xi}}\right)$
$\stackrel{{\scriptstyle{\mathbb{P}}}}{{\simeq}}{\mathop{\mathbb{E}}}_{\bm{Z}^{0},\dots,\bm{Z}^{t},\bm{Z}_{\bm{W}_{0}}}\left[\Phi_{N}\left(\bm{Z}_{\bm{W}_{0}}{\bm{\rho}}_{\bm{W}_{0}}^{-1}\bm{\nu}^{0}+\bm{W}_{0}\hat{\bm{\nu}}^{0}+\bm{Z}^{0},\dots,\bm{Z}_{\bm{W}_{0}}{\bm{\rho}}_{\bm{W}_{0}}^{-1}\bm{\nu}^{t}+\bm{W}_{0}\hat{\bm{\nu}}^{t}+\bm{Z}^{t},\bm{Z}_{\bm{W}_{0}};\bm{W}_{0},{\bm{\Xi}}\right)\right].$ (44)
@@
Proof  The main differences between the AMP result in Gerbelot and Berthier (2023) and our characterization are: 1) the $\varphi$ function is allowed to depend on an auxiliary random variable ${\bm{\Xi}}$, 2) the test function $\Phi_{N}$ is allowed to depend additionally on $\bm{A}\bm{W}_{0}$, and on the fixed parameters $\bm{W}_{0}$ and ${\bm{\Xi}}$. Assumption (B2) guarantees that $\tilde{f}^{t}$ maintains the same required convergence properties despite modification 1)1)1 ).
We now address modification 2). Since ${\bm{\Xi}}$ and $\bm{W}_{0}$ are fixed and $\tilde{\Phi}_{N}:(\bm{X}^{0},\dots,\bm{X}^{t},\bm{V})\mapsto\Phi_{N}(\mathbf{X}^{0},\mathbf{X}^{1},...,\mathbf{X}^{t},\bm{V};\bm{W}_{0},{\bm{\Xi}})$ is assumed to be uniformly pseudo-Lipschitz, these can be included in $\Phi_{N}$ and the right hand side of (44) is unaffected. Further, note that by Proposition 2 in Gerbelot and Berthier (2023) (a standard upper bound on the operator norm of $\bm{A}\mathbin{\overset{}{\kern 0.0pt\sim}}GOE(N)$) we have that $\|\bm{A}\bm{W}_{0}\|_{F}/\sqrt{N}<\infty$ almost surely as $N\to\infty$. We can therefore include $\bm{A}\bm{W}_{0}$ in the definition of $\Phi_{N}$ as well, and the uniform pseudo-Lipschitzness of $\tilde{\Phi}_{N}$ guarantees convergence to the result.
@@
for any $\bm{W}\in{\mathbb{R}}^{N\times L}$. We therefore have $\varphi(\bm{A}\bm{W}_{0},{\bm{\Xi}})=q\left(\bm{X}\bm{B},{\bm{\Psi}},\bm{\varepsilon}\right)$. Let $\tilde{f}^{t}:{\mathbb{R}}^{N\times L}\times{\mathbb{R}}^{N}\to{\mathbb{R}}^{N\times L}$ such that
$\begin{split}&\tilde{f}^{2t+1}(\varphi(\bm{A}\bm{W}_{0},{\bm{\Xi}}),\bm{U})=\sqrt{\frac{\delta+1}{\delta}}\begin{bmatrix}g^{t}(\bm{U}_{[n]},q(\bm{X}\bm{B},{\bm{\Psi}},\bm{\varepsilon}))\\ {\bm{0}}_{p\times L}\end{bmatrix}\quad\text{and}\\ &\tilde{f}^{2t}(\varphi(\bm{A}\bm{W}_{0},{\bm{\Xi}}),\bm{U})=\sqrt{\frac{\delta+1}{\delta}}\begin{bmatrix}\bm{0}_{n\times L}\\ f^{t}(\bm{U}_{[-p]})\end{bmatrix},\end{split}$ (47)
@@
$\bm{X}^{2t+1}$ $=\begin{bmatrix}\bm{X}f^{t}(\bm{X}^{2t}_{[-p]})-g^{t-1}(\bm{X}^{2t-1}_{[n]},q(\bm{X}\bm{B},{\bm{\Xi}}))\cdot\overbrace{\frac{1}{n}\left(\sum_{i=1}^{p}\frac{\partial{f^{t}(\bm{X}^{2t}_{[-p]})_{i}}}{\partial{\bm{X}^{2t}_{[-p],i}}}\right)^{\top}}^{=(\bm{F}^{t})^{\top}}\\ \bm{D}_{2}f^{t}(\bm{X}^{2t}_{[-p]})\end{bmatrix},$ (48)
$\bm{X}^{2t}$ $=\begin{bmatrix}\bm{D}_{1}g^{t-1}(\bm{X}^{2t-1}_{[n]},q(\bm{X}\bm{B},{\bm{\Xi}}))\\ \bm{X}^{\top}g^{t-1}(\bm{X}^{2t-1}_{[n]},q(\bm{X}\bm{B},{\bm{\Xi}}))-f^{t-1}(\bm{X}^{2t-2}_{[-p]})\cdot\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\partial{g^{t-1}(\bm{X}^{2t-1}_{[n]},q(\bm{X}\bm{B},{\bm{\Xi}}))}}{\partial{\bm{X}^{2t-1}_{[n],i}}}\right)^{\top}}_{=(\bm{C}^{t-1})^{\top}}\end{bmatrix},$ (49)

### v3

@@ 4.2 Rectified Linear Regression with Change Points
For the rectified linear model in (4) with a single change point at $(0.6n)$, Figure 5 (right) plots the Hausdorff distance divided by $n$ as a function of $\delta$. We let $p=600$, $L=3$, ${\mathbb{P}}_{\bar{\bm{B}}}={\sf N}({\bm{0}},\delta\bm{I})$, $\sigma=0.3$, and average over ten trials. Similarly, Table 1 displays the performance of AMP in estimating the number of change points in the same setup, where we vary $\sigma$ from $0.5$ to $0.3$. We observe a close match between the AMP performance and that dictated by theory in Proposition 3. Figure 2 (right) plots the estimated posterior $p_{\bar{{\bm{\Psi}}}|\bm{V}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}(\cdot|\bm{\Theta}^{t},\bm{y})$ against its the limiting posterior $p_{\bar{{\bm{\Psi}}}|\bm{V}_{\bm{\Theta}}^{t},q(\bm{Z},\bar{{\bm{\Psi}}},\bar{\bm{\varepsilon}})}(\cdot|{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},{\bm{\Psi}},\bm{\varepsilon}))$, for the rectified linear model with $p=600,\sigma=0.5$, true change point locations at $n/3$ and $8n/15$, and $L=3$ with a uniform prior over all configurations with two change points. We empirically observe a match between the estimated and limiting posteriors as per Proposition 4.
@@ 4.3 Logistic Regression with Change Points
[…]
@@
Figure 2 (middle) shows the estimated posterior $p(\cdot|\bm{\Theta}^{t},\bm{y})$ from (34) and the limiting (exact) posterior $p(\cdot|{\bm{V}}_{\bm{\Theta}}^{t},q(\bm{Z},{\bm{\Psi}},\bm{\varepsilon}))$ in the logistic model with two true change points at $n/3$ and $8n/15$. We let $p=600$, ${\mathbb{P}}_{\bar{\bm{B}}}={\sf N}({\bm{0}},20\bm{I})$, and subsample the domain $[0,1]$ by a factor of eight. We let $L=3$ and place uniform prior over all possible two-changepoint configurations. We observe a close match with the theory in Proposition 4, and concentration around the true change point locations as $\delta$ increases.
Figure 6 compares AMP with the binary segmentation algorithm proposed in Wang et al. (2023) for sparse logistic regression. We use $p=300,\Delta=0.15n$, and a sparse Bernoulli-Gaussian signal prior ${\mathbb{P}}_{\bar{\bm{B}}}=0.5{\sf N}({\bm{0}},15\delta\bm{I})+0.5\delta_{\bm{0}}$. The true change points are at $n/3$ and $8n/15$ with $L^{*}=3$. The hyperparameters for binary segmentation are chosen as suggested in the source code of Wang et al. (2023). Both algorithms can take advantage of prior information on the number of change points. In Fig. 6a, both algorithms assume a maximum of $L=3$ or 4 signals while $L^{*}=3$. The algorithms are observed to achieve similar Hausdorff distance, with AMP enjoying a much lower runtime when assuming $L=3$. Fig. 6b compares the two algorithms by assuming there are exactly 3 signals. The performance of AMP improves significantly with larger $\delta$, both in terms of Hausdorff distance and runtime. We recall that the algorithm in Wang et al. (2023) is tailored to sparse logistic regression, whereas AMP can account for a broader range of signal priors.
@@
(a) Algorithms assume a maximum of $3$ or 44 signals (that is, $L=3$ or 4).
@@
[…]
@@
5 Discussion
In this work we proposed an AMP algorithm for regression with change points. The algorithm can efficiently incorporate prior information on the change points, and its state evolution characterization gives rigorous asymptotic guarantees on estimation accuracy. It also allows to quantify uncertainty in the change point estimates via a computable posterior distribution.
In principle, the choice of change point prior could significantly affect the performance of our proposed algorithm—for example, the performance would be poor if we choose a prior that assigns very small probability mass to the true change-point locations. However, our experimental results indicate that an uninformative prior (uniform on both the number and locations of the change points) consistently gives good performance across a range of settings. The algorithm also uses priors on the signal $\bm{B}$ and the measurement noise $\bm{\varepsilon}$ to tailor the AMP denoising functions $f^{t},g^{t}$. A misspecified prior on $\bm{\varepsilon}$ or $\bm{B}$ would lead to choices on $f^{t},g^{t}$ that are suboptimal, but we emphasize the algorithm and the theoretical guarantees remain valid: the performance of the algorithm is characterized by state evolution parameters, which in turn depend on the denoising functions used within AMP. Figures 3, 4, 6, 7, 8 illustrate the robustness of our algorithm to lack of exact knowledge about signal or noise priors.
A limitation of our method is that the AMP algorithm and its theoretical guarantees are derived assuming that the covariates are i.i.d. Gaussian. We discuss below how this assumption can be relaxed in certain settings, along with other directions for future work.
Non-Gaussian Design
Based on recent universality results for AMP (Wang et al., 2024), we expect that our algorithm and theoretical guarantees will apply to a broad class of i.i.d. designs. We also expect that our approach can be generalized to the much broader class of rotationally invariant designs, for which AMP-like algorithms have been proposed for regression without change points (Ma and Ping, 2017; Rangan et al., 2019; Takeuchi, 2020; Pandit et al., 2020).
Gaussian Design with General Covariance
In the setting where $\bm{X}_{i}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\sf N}({\bm{0}},{\bm{\Sigma}})$ for $i\in[n]$ and for some covariance matrix ${\bm{\Sigma}}\in{\mathbb{R}}^{p\times p}$, we can write $\bm{X}_{i}={\bm{\Sigma}}^{1/2}\tilde{\bm{X}}_{i}$, where $\tilde{\bm{X}}_{i}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\sf N}({\bm{0}},\bm{I}_{p})$ for $i\in[n]$. The model (1) can then be equivalently written as:
$y_{i}=q\big{(}(\tilde{\bm{X}}_{i})^{\top}({\bm{\Sigma}}^{1/2}{\bm{\beta}}^{(i)}),\,\varepsilon_{i}\big{)},\quad i=1,\dots,n,$
where $\tilde{\bm{X}}\in{\mathbb{R}}^{n\times p}$ is now our effective design matrix composed of stacked rows $(\tilde{\bm{X}}_{i})^{\top}$, and ${\bm{\Sigma}}^{1/2}{\bm{\beta}}^{(i)}$ is our effective signal. If ${\bm{\Sigma}}\in{\mathbb{R}}^{p\times p}$ is known or can be approximated with reasonable accuracy, the user can run the AMP iteration (LABEL:eq:amp) with $\tilde{\bm{X}}$ in place of $\bm{X}$. Since the entries of $\tilde{\bm{X}}$ are i.i.d. Gaussian, we can apply Theorem 1 to this iteration to obtain guarantees on the accuracy of change point estimation (via Proposition 3) and guarantees on the recovery of ${\bm{\Sigma}}^{1/2}\bm{B}$ (via Theorem 1). Because $\bm{\Sigma}$} is known, this also gives us guarantees on the recovery of $\bm{B}$.
The more realistic setting of unknown $\bm{\Sigma}$} is challenging, and developing effective AMP-like algorithms for this case is an interesting open question. One idea is to use weighted versions of standard convex estimators for high-dimensional regression such as LASSO or M-estimators, where the weights can be chosen to encode prior information on change point locations. By leveraging the connection between AMP and convex optimization (Section 4.4 of Feng et al. 2022), it is possible to analyze such convex estimators under unknown covariance $\bm{\Sigma}$}, as in (Zhao et al., 2022).
Design with Temporal Dependence
Xu et al. (2024) recently studied the problem of change point detection in linear regression with temporally dependent data, where the rows of the design matrix $\bm{X}$ form a high-dimensional time-series. An AMP framework has recently been proposed for linear regression when the data are generated from a high-dimensional time series (Tieplova et al., 2025). Using this framework to generalize our change point AMP to temporal dependent data is a promising direction for future work.
@@ Appendix A Proof of State Evolution
We extend this lemma by considering an independent random matrix ${\bm{\Xi}}\in{\mathbb{R}}^{N\times L_{{\bm{\Xi}}}}$, serving as input to a set of pseudo-Lipschitz functions $\varphi:\mathbb{R}^{N\times(Q+L_{{\bm{\Xi}}})}\to\mathbb{R}^{N}$. We then analyze a similar AMP iteration whose denoising function $\tilde{f}^{t}$ takes $\varphi(\bm{A}\bm{W}_{0},{\bm{\Xi}})$ as input instead of $\varphi(\bm{A}\bm{W}_{0})$, initialized with an independent initializer $\bm{X}^{0}\in\mathbb{R}^{N\times Q}$ and $\bm{M}^{-1}:={\bm{0}}\in{\mathbb{R}}^{N\times Q}$:
@@
    $\bm{A}\in\mathbb{R}^{N\times N}$ is a GOE($N$) matrix, that is, $\bm{A}=\mathbf{G}+\mathbf{G}^{\top}$ for $\mathbf{G}\in\mathbb{R}^{N\times N}$ with i.i.d. entries $G_{ij}\sim{\sf N}(0,1/(2N))$.
@@
    where $(\bm{Z}^{s},\bm{Z}^{t})\in(\mathbb{R}^{N\times Q})^{2}$$(\bm{Z}^{s},\bm{Z}^{t})\sim{\sf N}(0,\bm{\kappa}\otimes\mathbf{I}_{N})$.
@@
$\Phi_{N}\left(\bm{X}^{0},\bm{X}^{1},...,\bm{X}^{t},\bm{A}\bm{W}_{0};\bm{W}_{0},{\bm{\Xi}}\right)\stackrel{{\scriptstyle{\mathbb{P}}}}{{\simeq}}$
${\mathop{\mathbb{E}}}_{\bm{Z}^{1},\dots,\bm{Z}^{t},\bm{Z}_{\bm{W}_{0}}}\left[\Phi_{N}\bigl{(}\bm{X}^{0},\bm{Z}_{\bm{W}_{0}}{\bm{\rho}}^{-1}_{\bm{W}_{0}}{\bm{\nu}}^{1}+\bm{W}_{0}\hat{{\bm{\nu}}}^{t}+\bm{Z}^{1},\dots,\right$.
$\hskip 142.26378pt\left.\bm{Z}_{\bm{W}_{0}}{\bm{\rho}}_{\bm{W}_{0}}^{-1}\bm{\nu}^{t}+\bm{W}_{0}\hat{\bm{\nu}}^{t}+\bm{Z}^{t},\bm{Z}_{\bm{W}_{0}};\bm{W}_{0},{\bm{\Xi}}\bigr{)}\right].$ (44)
@@
Proof  The main differences between the AMP result in Gerbelot and Berthier (2023) and our characterization are: 1) the $\varphi$ function is allowed to depend on an auxiliary random variable $\bm{\Xi}$}, 2) the test function $\Phi_{N}$ is allowed to depend additionally on $\bm{A}\bm{W}_{0}$, and on the fixed parameters $\bm{W}_{0}$ and $\bm{\Xi}$}. Assumption (B2) guarantees that $\tilde{f}^{t}$ maintains the same required convergence properties despite modification $1)$.
We now address modification 2). Since $\bm{\Xi}$} and $\bm{W}_{0}$ are fixed and $(\bm{X}^{0},\dots,\bm{X}^{t},\bm{V})\mapsto\Phi_{N}(\mathbf{X}^{0},\mathbf{X}^{1},...,\mathbf{X}^{t},\bm{V};\bm{W}_{0},{\bm{\Xi}})$ is assumed to be uniformly pseudo-Lipschitz, these can be included in $\Phi_{N}$ and the right hand side of (44) is unaffected. We now present the argument for why $\bm{A}\bm{W}_{0}$ can be included as in input to $\Phi_{N}$ on the LHS of (44), and yield a corresponding input $\bm{Z}_{\bm{W}_{0}}$ on the RHS of (44). Consider the following “augmented” AMP iteration, initialized with $\mathbf{X}^{aug,0}:=\begin{bmatrix}\bm{X}^{0}&\bm{X}^{0}\end{bmatrix}\in\mathbb{R}^{N\times 2Q}$ and $\bm{M}^{aug,-1}:={\bm{0}}\in{\mathbb{R}}^{N\times 2Q}$:
$\bm{X}^{aug,t+1}=\bm{A}\bm{M}^{aug,t}-\bm{M}^{aug,t-1}(\bm{b}^{aug,t})^{\top}$ $\in\mathbb{R}^{N\times 2Q}\,$, (45)
$\bm{M}^{aug,t}=\tilde{f}^{aug,t}(\varphi^{aug}\left(\bm{A}\bm{W}^{aug}_{0},{\bm{\Xi}}\right),\bm{X}^{aug,t})$ $\in\mathbb{R}^{N\times 2Q}\,$, (46)
$\bm{b}^{aug,t}=\frac{1}{N}\sum_{i=1}^{N}\frac{\partial\tilde{f}^{aug,t}_{i}}{\partial\bm{X}_{i}^{aug,t}}(\varphi^{aug}\left(\bm{A}\bm{W}^{aug}_{0},{\bm{\Xi}}\right),\bm{X}^{aug,t})$ $\in\mathbb{R}^{2Q\times 2Q}\,.$ (47)
In the above iteration, we define $\bm{W}_{0}^{aug}:=\begin{bmatrix}\bm{W}_{0}&\bm{W}_{0}\end{bmatrix}\in{\mathbb{R}}^{N\times 2Q},\varphi^{aug}\left(\bm{A}\bm{W}^{aug}_{0},{\bm{\Xi}}\right):=\begin{bmatrix}\varphi\left(\bm{A}\bm{W}_{0},{\bm{\Xi}}\right)&\varphi\left(\bm{A}\bm{W}_{0},{\bm{\Xi}}\right)\end{bmatrix}\in{\mathbb{R}}^{N\times 2Q}$. Letting $\bm{X}^{t+1},\tilde{\bm{X}}^{t+1}\in{\mathbb{R}}^{N\times Q}$ take on the definition
$\begin{bmatrix}\bm{X}^{t+1}&\tilde{\bm{X}}^{t+1}\end{bmatrix}:=\bm{X}^{aug,t+1}$
and defining:
$\tilde{f}^{aug,t}(\varphi^{aug}\left(\bm{A}\bm{W}^{aug}_{0},{\bm{\Xi}}\right),\bm{X}^{aug,t}):=\begin{bmatrix}\tilde{f}^{t}(\varphi\left(\bm{A}\bm{W}_{0},{\bm{\Xi}}\right),\bm{X}^{t})&\xi^{t}(\tilde{\bm{X}}^{t})\end{bmatrix}$
for some uniformly pseudo-Lipschitz function $\xi^{t}(\tilde{\bm{X}}^{t+1})$ with $\xi^{0}(\tilde{\bm{X}}^{0}):=\bm{W}_{0}$, we obtain that:
$\bm{M}^{aug,t}=:\begin{bmatrix}\bm{M}^{t}&\tilde{\bm{M}}^{t}\end{bmatrix}$ (48)
$\bm{b}^{aug,t}=\begin{bmatrix}\bm{b}^{t}&{\bm{0}}\\ {\bm{0}}&*\end{bmatrix}$, (49)
where $\bm{M}^{t}$ and $\bm{b}^{t}$ are quantities from the initial iteration (36)–(38), and * denotes an unspecified matrix in ${\mathbb{R}}^{Q\times Q}$. We therefore have that the first $Q$ columns of the matrix iteration (45)–(47) are equivalent to the original iteration (36)–(38). The last $Q$ columns of the augmented iteration will be used to incorporate $\bm{A}\bm{W}_{0}$.
For $t\geq 1$, let $(\bm{X}^{aug,t})_{[:,:Q]},(\bm{X}^{aug,t})_{[:,Q:]}$ denote the $N\times Q$ matrices composed of the first $Q$ columns and of the last $Q$ columns of $\bm{X}^{aug,t}$, respectively. Let ${\bm{\rho}}:=\frac{1}{N}\bm{W}_{0}^{\top}\bm{W}_{0}$, and for $t\geq 1$ define the following random variables, independent of all else: $(\bm{Z}_{\bm{W}_{0}})_{i}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\sf N}({\bm{0}},{\bm{\rho}})$ and $(\bm{Z}^{t})_{i}\mathbin{\overset{i.i.d}{\kern 0.0pt\sim}}{\sf N}(0,{\bm{\kappa}}^{t,t})$ for $i\in[n]$. We compute the state evolution equations associated with the augmented AMP iteration for $t=1$, outlined in (C.25)(C.28)(C.25)-(C.28) of Gerbelot and Berthier (2023), to obtain:
${\bm{\nu}}^{aug,1}=\lim_{N\to\infty}\frac{1}{N}{\mathop{\mathbb{E}}}\left[\begin{bmatrix}\bm{W}_{0}^{\top}\\ \bm{W}_{0}^{\top}\end{bmatrix}\begin{bmatrix}\tilde{f}^{0}(\bm{X}^{0})&\xi^{0}(\tilde{\bm{X}}^{0})\end{bmatrix}\right]=\begin{bmatrix}{\bm{\nu}}^{1}&{\bm{\rho}}\\ {\bm{\nu}}^{1}&{\bm{\rho}}\end{bmatrix}$,
$\hat{{\bm{\nu}}}^{aug,1}=\begin{bmatrix}\hat{{\bm{\nu}}}^{1}&{\bm{0}}\\ {\bm{0}}&{\bm{0}}\end{bmatrix}$,
${\bm{\kappa}}^{aug,1,1}=\lim_{N\to\infty}\frac{1}{N}$
[…]
$\hskip 35.56593pt=\begin{bmatrix}{\bm{\kappa}}^{1,1}&{\bm{0}}\\ {\bm{0}}&{\bm{0}}\end{bmatrix}$.
Applying Lemma 14 in Gerbelot and Berthier (2023) to the pseudo-Lipschitz function
$\tilde{\Phi}_{N}\left(\bm{X}^{aug,1},\dots,\bm{X}^{aug,t}\right):=\Phi_{N}\left((\bm{X}^{aug,1})_{[:,:Q]},\dots,(\bm{X}^{aug,t})_{[:,:Q]},(\bm{X}^{aug,1})_{[:,Q:]}\right),$
we obtain:
$\Phi_{N}\left(\bm{X}^{1},\dots,\bm{X}^{t},\bm{A}\bm{W}_{0}\right)$
$\Phi_{N}\left((\bm{X}^{aug,1})_{[:,:Q]},\dots,(\bm{X}^{aug,t})_{[:,:Q]},(\bm{X}^{aug,1})_{[:,Q:]}\right)$
$=\tilde{\Phi}_{N}\left(\bm{X}^{aug,1},\dots,\bm{X}^{aug,t}\right)$
[…]
$={\mathop{\mathbb{E}}}\tilde{\Phi}_{N}\left(\begin{bmatrix}\bm{Z}_{\bm{W}_{0}}{\bm{\rho}}^{-1}{\bm{\nu}}^{1}+\bm{W}_{0}\hat{{\bm{\nu}}}^{1}+\bm{Z}^{1}&\bm{Z}_{\bm{W}_{0}}\end{bmatrix}$$\dots,\right$.
[…]
$={\mathop{\mathbb{E}}}\Phi_{N}\left(\bm{Z}_{\bm{W}_{0}}{\bm{\rho}}^{-1}{\bm{\nu}}^{1}+\bm{W}_{0}\hat{{\bm{\nu}}}^{1}+\bm{Z}^{1},\cdots,\bm{Z}_{\bm{W}_{0}}{\bm{\rho}}^{-1}{\bm{\nu}}^{t}+\bm{W}_{0}\hat{{\bm{\nu}}}^{t}+\bm{Z}^{t},\bm{Z}_{\bm{W}_{0}}\right),$
and the result follows.
@@
for any $\bm{W}\in{\mathbb{R}}^{N\times L}$. We therefore have $\varphi(\bm{A}\bm{W}_{0},{\bm{\Xi}})=q\left(\bm{X}\bm{B},{\bm{\Psi}},\bm{\varepsilon}\right)$. Let
$\tilde{f}^{t}:{\mathbb{R}}^{N\times L}\times{\mathbb{R}}^{N}\to{\mathbb{R}}^{N\times L}$
such that
[…]
@@
$\bm{X}^{2t+1}$
$=\begin{bmatrix}\bm{X}f^{t}(\bm{X}^{2t}_{[-p]})-g^{t-1}(\bm{X}^{2t-1}_{[n]},q(\bm{X}\bm{B},{\bm{\Xi}}))\cdot\overbrace{\frac{1}{n}\left(\sum_{i=1}^{p}\frac{\partial{f^{t}(\bm{X}^{2t}_{[-p]})_{i}}}{\partial{\bm{X}^{2t}_{[-p],i}}}\right)^{\top}}^{=(\bm{F}^{t})^{\top}}\\ \bm{D}_{2}f^{t}(\bm{X}^{2t}_{[-p]})\end{bmatrix}$, (53)
$\bm{X}^{2t}$
$=\begin{bmatrix}\bm{D}_{1}g^{t-1}(\bm{X}^{2t-1}_{[n]},q(\bm{X}\bm{B},{\bm{\Xi}}))\\ \bm{X}^{\top}g^{t-1}(\bm{X}^{2t-1}_{[n]},q(\bm{X}\bm{B},{\bm{\Xi}}))-f^{t-1}(\bm{X}^{2t-2}_{[-p]})\cdot\underbrace{\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\partial{g^{t-1}(\bm{X}^{2t-1}_{[n]},q(\bm{X}\bm{B},{\bm{\Xi}}))}}{\partial{\bm{X}^{2t-1}_{[n],i}}}\right)^{\top}}_{=(\bm{C}^{t-1})^{\top}}\end{bmatrix}$, (54)